OpenStudy (anonymous):

For questions 6 – 9 prove the following trigonometric identities. 7. (sec x + tan x)((1 – sin x) / cos x) = 1 Calculation Reason

OpenStudy (sidsiddhartha):

take cosx/i-sinx multiply conjugate of the denominator (1+sin x) / (1+sin x) so the eqtn becomes cos x / (1 - sin x) * (1+sin x) / (1+sin x) =cos x (1 + sin x) / (1 - sin² x) =cos x (1 + sin x) / cos² x =(1 + sin x) / cos x =1/cos x + sin x / cos x =sec x + tan x hope helped

OpenStudy (anonymous):

thanks @sidsiddhartha

OpenStudy (anonymous):

now if i had some reasoning then you would be the best person ever @anyone